OpenStudy (anonymous):

what is the domain: f(x)=Square root x^4-81

OpenStudy (anonymous):

\[\sqrt{x ^{4}-81}\]

OpenStudy (anonymous):

\[(-\infty,-3), (3,\infty)\]

OpenStudy (anonymous):

thanks